OpenStudy (anonymous):

differentiate y=tan2 (x/3) help me...

OpenStudy (anonymous):

\[y = \tan^{2} (\frac{ x }{ 3 })\]

OpenStudy (science0229):

You have to use the chain rule twice.

OpenStudy (science0229):

First, let x/3=u

OpenStudy (science0229):

\[y=\tan^2u\]

OpenStudy (science0229):

Differentiate it like the following;\[\frac{ dy }{ dx }=\frac{ dy }{ du }*\frac{ du }{ dx }\]

OpenStudy (science0229):

\[\frac{ dy }{ du }=2\tan u \sec^2u\]

OpenStudy (science0229):

\[\frac{ du }{ dx }=\frac{ 1 }{ 3 }\]

OpenStudy (science0229):

Therefore, \[\frac{ dy }{ dx }=\frac{ 2 }{ 3 }\tan u \sec^2 u=\frac{ 2 }{ 3 }\tan \frac{ x }{ 3 } \sec^2 \frac{ x }{ 3 }\]

OpenStudy (science0229):

Oh, and my bad. You only have to use the chain rule once, not twice.

OpenStudy (science0229):

Any questions?

OpenStudy (anonymous):

okay!! thank u very much.